### v1

@@ 3.2 GitHub Data Collection and Prepossessing
We utilized the same SWSs listed in Table 1 for our analysis on GitHub. To gather data on issues and pull requests, we employed the GitHub REST API [133]. We observed that several SWSs are distributed across multiple repositories during this process. For instance, the Galaxy workflow system spans several repositories (e.g., tools-iuc, planemo, galaxy-hub) related to its ecosystem. Similarly, Nextflow is associated with multiple repositories (e.g., patterns, rnaseq-nf, nf-validation). Consequently, we aggregated all issues and pull requests from these relevant repositories. We treated each issue report and pull request as an individual document in the topic model. We joined the title and body of each issue and pull requests as a single document. We remove code snippets, URLs, and images from the body. Additionally, we eliminate stop words, non-alphabetical characters, and punctuation to refine the textual data. As of January 31, 2024, after preprocessing, our dataset comprises 140,400 issues and pull requests collected from various GitHub repositories. In our replication package [51], we provided the GitHub link to the corresponding SWS, along with the data collection scripts and the entire dataset.
@@
In BERTopic, there is a parameter nr_topics to control the number of topics, which merges topics after they have been created by supporting the creation of a fixed number of topics. However, it is recommended that the number of topics be controlled through a cluster model. We utilized the HDBSCAN [135] cluster model to perform this task. We improved the default representation of topics using the Countvectorizer  [^2] https://scikit-learn.org/stable/modules/generated/sklearn.feature_extraction.text.CountVectorizer.html. It helps to ignore infrequent words and increase the n-gram range. Following previous works [21, 17], we used the unigram and bigram models. Finally, we obtained 10 topics for SO data and 13 topics from the GitHub data. BERTopic includes a -1 topic to group documents that do not align well with the identified topics. This situation may arise when a document contains excessive noise, such as stopwords or irrelevant information, or does not strongly connect to the main themes identified by the model. The -1 topic serves as a ’catch-all’ for these outlier documents. However, since we preprocessed the data to eliminate such elements before applying BERTopic, our main topic began with the label -1. The scripts for data preprocessing and topic generation for SO and GitHub data are provided in the replication package [51]. We described our findings in the following Section.
@@ 4 Results
In this section, we present the analysis of the SWS-related data we obtained from SO and GitHub and the topics to answer our research questions.
@@
Motivation: Scientific Workflow Systems (SWSs) development has unique characteristics that set it apart from traditional software development [59]. For example, SWSs developers need expertise in complex data management and the involvement of specialized software, high-performance resources, and hardware. SWSs should be adaptable to frequent changes, and adherence to community standards, reproducibility, and computational efficiency are crucial. Multidisciplinary team involvement and usage of domain specific languages are often necessary. These particularities make SWS development a complex endeavor, requiring a deep understanding of scientific domains and advanced software engineering practices, which are often unnecessary or overlooked in most conventional software development tasks. Hence, the challenges SWSs developers face will likely differ from those of traditional software development. Developers use Q&A websites to communicate both problems and solutions, and they rely on GitHub for version control and collaboration, including code review, project management, and collective software development. This research question aims to analyze invaluable data from SO and GitHub to identify the most common and pressing topics related to SWSs and to explore the relationships between these topics on both platforms. By pinpointing the challenges frequently encountered by the SWSs community, we can highlight which topics are gaining traction and which are proving challenging to address. Understanding these trends is crucial for supporting the SWSs community and guiding future development efforts.
Approach: We used BERTopic to identify the different topics developers discuss on SO and GitHub, as mentioned in Section 3. The first three authors (annotators) labeled the set of topics based on the overall theme of the posts. Each annotator individually inspected the top 15 keywords and a random sample of at least 30 posts from each topic to assign a title that best represents the content. The authors then discussed the labels for each of the 10 topics to reach a consensus on their titles. We followed a similar approach to label topics for the GitHub issues and pull requests. Additionally, we examined the most popular SWSs topics among developers. To investigate this, we used two complementary measurements of popularity that have been adopted in prior works [17, 109, 19, 136, 118]:
@@
Results Obtained Using SO: Following the process described in sub-section 3.3, we identified 10 topics for SO data. Table 2 presents the obtained ten topic titles along with their associated keywords. It also shows the number of posts related to each topic and the popularity of these topics based on our metrics: views and scores received by developers on SO. As shown in Table 2, developers inquire about various topics in SWSs development, with the number of posts varying across topics. Below, we discuss these topics in greater detail.
@@
3. Processor for Data Processing: SWSs provide a rich set of processors covering a wide range of data integration, transformation, routing, and interaction tasks. Practitioners can design sophisticated data pipelines to meet their specific data processing needs by selecting and configuring the appropriate processors. This topic covers 14.02% of the posts. One example of this topic with about 27K views is ’How to use NiFi ExecuteScript processor with Python?’. ExecuteScript processor allows users to write and execute custom scripts within a data flow. This processor provides flexibility for handling complex processing logic that the built-in processors do not cover. Here, the developer tried to use the ExecuteScript processor with a simple Python script but failed to see the output and asked for help to explain how to use it with an example. Another example of this topic with 12K views is ’How to specify output filename with PutFile processor?’. Here, the problem was that the PutFile processor was saving files using a timestamp, but the expectation was to save them with the original filename. Thus, the developer is seeking a solution. This topic indicates developers’ challenges in understanding data processing processors.
@@
8. Data Transformation: In SWS, this topic discusses converting data from one format or structure to another, enabling easier analysis, visualization, or integration with other datasets. This involves altering the shape or appearance of data objects while preserving their content and, for instance, using libraries like Jolt to transform JSON data from one format to another. This topic covers around 1.26% of posts of our SO dataset. One example of this topic, with around 10K views, is titled ’Is it possible to concatenate the values of JSON attributes using JOLT?’. In this instance, the developer inquired whether JOLT could be used to concatenate the values of three JSON attributes into a new attribute and also asked potential alternative approaches.

### v2

@@ 3.2 GitHub Data Collection and Preprocessing
Step 1: Collecting GitHub Issues and Pull Requests. We utilized the same SWSs listed in Table 1 for our analysis on GitHub. To gather data on issues and pull requests, we employed the GitHub REST API [137]. We observed that several SWSs are distributed across multiple repositories during this process. For instance, the Galaxy workflow system spans several repositories (e.g., tools-iuc, planemo, galaxy-hub) related to its ecosystem. Similarly, Nextflow is associated with multiple repositories (e.g., patterns, rnaseq-nf, nf-validation). Consequently, we aggregated all issues and pull requests from these relevant repositories.
As of January 31, 2024, our initial GitHub dataset comprised 169,692 issues and pull requests. However, we observed that many pull requests were directly linked to issues, and treating them as independent entries could introduce redundancy into the analysis. To address this, we utilized the GitHub Search API to identify and exclude 6,574 PRs that were associated with existing issues. After this refinement, our final dataset consisted of 163,118 distinct and non-redundant issues and PRs. The distribution of this data is presented in Table 1. All scripts used for data collection, PR-issue linkage detection, and filtering are provided in our replication package [53] to ensure transparency and reproducibility.
Step 2: Preprocessing GitHub Data. To prepare the GitHub issues and pull requests for topic modeling, we applied a series of preprocessing steps to clean and normalize the textual data. First, we concatenated the title and body of each issue and pull request to form a single document, ensuring sufficient context for topic extraction. We then removed code snippets enclosed in backticks using regular expressions, as such blocks often contain syntax elements irrelevant to semantic modeling. Next, we eliminated URLs, which typically reference external documentation or repositories, and cleaned out any remaining HTML tags using regular expressions. To further reduce noise, we removed special characters and punctuation. We also applied stopword removal using the NLTK stopword list and supplemented it by filtering out project-specific terms such as airflow, celery, and nextflow to prevent these frequently occurring system names from dominating the topic space. Finally, we performed lemmatization using spaCy’s en_core_web_sm model to convert words into their base forms, thus improving topic coherence and reducing lexical variation.
Removing Code Snippets: To remove code snippets from GitHub issues and pull requests, we used regular expressions to detect Markdown-style formatting. We identified and removed inline code enclosed within single backticks and multi-line code blocks enclosed in triple backticks. This preprocessing step was essential to prevent programming syntax and formatting elements from distorting the semantic structure of the text during topic modeling.
def remove_code_snippets(text):
if not text:
return text
# Remove inline code (e.g., ‘code‘)
text = re.sub(r’‘[^‘]*‘’, ’’, text)
# Remove block code (e.g., ‘‘‘code‘‘‘)
[…]
return text
Remove URLs: Many issues/PRs contain links to external documentation, repositories, or Stack Overflow discussions. These URLs do not contribute to topic modeling; thus, we removed them using regular expressions.
def remove_https_links(text):
if isinstance(text, str):
# Regular expression pattern to match HTTPS links
https_pattern = r’https:/$\S$+’
# Remove all matches from the text
return re.sub(https_pattern, ’’, text).strip()
else:
# If not a string, return as is
return text
Removing HTML Tags: We removed HTML tags using regular expression.
def remove_html_tags(text):
if text is None:
return ’’
return re.sub(r’<.*?>’, ’’, text)
Removing Stopwords & Project-Specific Terms: We applied the same approach for removing stop words and project-specific terms as described in the SO data preprocessing step.
Removing Special Characters and Punctuation: We utilized the same approach as we did at SO data preprocessing for this step
Filtering Low-Information Issues and Pull Requests: After completing all the preprocessing steps, we identified a number of GitHub issues and pull requests containing only one or two words, such as update, Jpython, typo, or Add Tags, which lacked meaningful context or insight. To improve data quality, we removed entries where the text length was fewer than eight characters. This filtering step resulted in the exclusion of 636 issues and pull requests, yielding a final curated dataset of 162,169 GitHub issues and pull requests.
Lemmatization: For this step, we again follow the process we used in SO data lemmatization.
@@
In BERTopic, the nr_topics parameter allows control over the number of topics by merging similar ones after their initial creation. However, best practices recommend using a clustering model to determine the topic structure more naturally [26]. To enable accurate and coherent clustering, we first reduced the dimensionality of the document embeddings, as high-dimensional data can impair clustering performance. Among dimensionality reduction techniques, UMAP [139] is widely recognized for its ability to preserve local structure when projecting data to lower dimensions. Therefore, we employed UMAP for this step. We then applied HDBSCAN [140], a density-based clustering algorithm that complements UMAP well due to its sensitivity to local structure. Unlike traditional clustering methods, HDBSCAN can identify noise and does not force every data point into a cluster, making it especially effective for short, noisy texts such as SO posts and GitHub issues
For BERTopic modeling on both datasets, we carefully tuned UMAP parameters to balance topic granularity and semantic coherence. We experimented with n_neighbors values between 15 and 50 to ensure that semantically similar documents were grouped together while preserving a meaningful global topic structure. The n_components parameter was adjusted between 3 and 20 to retain essential semantic information from the high-dimensional sentence embeddings while enabling efficient clustering. We used cosine distance as the metric, as it aligns well with sentence-transformer embeddings and effectively captures angular similarity. For HDBSCAN, we tested min_cluster_size values from 50 to 300 in intervals of 10. We employed Euclidean distance for clustering, as it performs reliably in UMAP-reduced space, which is optimized for Euclidean geometry. Additionally, we used the eom method for cluster_selection_method because it identifies stable, well-separated clusters by emphasizing dense core regions. We improved the default representation of topics using the Countvectorizer  [^2] https://scikit-learn.org/stable/modules/generated/sklearn.feature_extraction.text.CountVectorizer.html. It helps to ignore infrequent words and increase the n-gram range. Following previous works [26, 22], we used the unigram and bigram models in Countvectorizer. The optimal topic coherence score of 0.64 was achieved for the GitHub dataset with n_neighbors = 20, n_components = 4, and min_cluster_size = 130, while the SO dataset yielded the optimal score of 0.63 with n_neighbors = 30, n_components = 3, and min_cluster_size = 210. This configuration enabled BERTopic to generate high-quality, interpretable topics while maintaining scalability across both dataset.
Using the above-mentioned parameters, we obtained 10 topics for SO data and 13 topics from the GitHub data. BERTopic includes a -1 topic to group documents that do not align well with the identified topics. This situation may arise when a document contains excessive noise, such as stopwords or irrelevant information, or does not strongly connect to the main themes identified by the model. The -1 topic serves as a ’catch-all’ for these outlier documents. However, since we preprocessed the data to eliminate such elements before applying BERTopic, our main topic began with the label -1. The scripts for data preprocessing and topic generation for SO and GitHub data are provided in the replication package [53]. We described our findings in the following Section.
@@ 4 Results
In this section, we presented the analysis of the SWS-related data we obtained from SO and GitHub and the topics to answer our research questions.
@@
Motivation: Scientific Workflow Systems (SWSs) development has unique characteristics that set it apart from traditional software development [61]. For example, SWSs developers need expertise in complex data management and the involvement of specialized software, high-performance resources, and hardware. SWSs should be adaptable to frequent changes, and adherence to community standards, reproducibility, and computational efficiency are crucial. Multidisciplinary team involvement and usage of domain specific languages are often necessary. These particularities make SWS development a complex endeavor, requiring a deep understanding of scientific domains and advanced software engineering practices, which are often unnecessary or overlooked in most conventional software development tasks. Hence, the challenges SWSs developers face may likely differ from those of traditional software development. Developers use Q&A websites to communicate both problems and solutions, and they rely on GitHub for version control and collaboration, including code review, project management, and collective software development. This research question aims to analyze invaluable data from SO and GitHub to identify the most common and pressing topics related to SWSs and to explore the relationships between these topics on both platforms. By pinpointing the challenges frequently encountered by the SWSs community, we can highlight which topics are gaining traction and which are proving challenging to address. Understanding these trends is crucial for supporting the SWSs community and guiding future development efforts.
Approach: To identify the key topics discussed by developers, we applied BERTopic to both the Stack Overflow and GitHub datasets, as described in Section 3. The BERTopic modeling process resulted in 10 distinct topics for the Stack Overflow data, 13 topics for the GitHub data, and the distribution of co-occurring words in each topic. We then manually assigned a meaningful label to each topic. Following established approaches in prior work [26, 141, 25, 24], the first author, with over four years of research experience in SWSs and more than nine years of professional software development experience, initially proposed candidate labels. These were derived from the top 15 keywords associated with each topic and a manual review of at least 30 random representative posts, issues, or pull requests per topic. Subsequently, the initial labels were reviewed by the first three authors of the paper in collaborative meetings. The second author contributes over eight years of expertise in SWS research, while the third author brings more than two decades of experience in empirical software engineering. Based on this iterative review and refinement process, we finalized meaningful and contextually accurate labels for all identified topics. Additionally, we examined the most popular SWSs topics among developers. To investigate this, we used two complementary measurements of popularity that have been adopted in prior works [22, 112, 24, 142, 121]:
@@
Results Obtained Using SO: Following the process described in sub-section 3.3, we identified 10 topics for SO data. Table 2 presents the obtained ten topic titles along with their associated keywords. It also shows the number of posts related to each topic and the popularity of these topics based on our metrics: views and scores received by developers on SO. As shown in Table 2, developers inquire about various topics in SWSs development, with the number of posts varying across topics. Below, we discussed these topics in greater detail.
@@
The challenges in Workflow Creation and Scheduling arose primarily from complexities in defining, managing, and executing Directed Acyclic Graphs (DAGs) in workflow engines like Apache Airflow. Users often struggle with dynamic job scheduling, ensuring task dependencies are correctly managed, and handling failures or retries efficiently. Issues related to rescheduling tasks, managing large-scale data workflows, and optimizing execution performance further complicate workflow automation. Additionally, debugging and maintaining DAGs, especially when integrating multiple processing steps, require a deep understanding of task execution order and error propagation, making workflow creation and scheduling a demanding task.
Table 3: Granular Subtopics under Workflow Creation and Scheduling
Subtopic Keywords Post Count Description Representative Example
DAG Syntax and Structural Issues dag, task, use, run, file, try, error, work, code 6116 Syntactic or structural issues in DAG definitions, including indentation, file format, and import errors. ’Run task using Python DAG; task not recognized due to indentation error in load definition.’
Cloud Composer Configuration and Integration composer, cloud, google, bigquery, use, cloud composer 1113 Cloud Composer-specific problems involving configuration, authentication, and deployment. ’Google Cloud Composer job fails, BigQuery hook not installed despite dependencies listed.’
DAG Scheduling and Triggering Problems dag, run, task, schedule, trigger, time 1013 DAG scheduling inconsistencies, task triggering issues, and incorrect start or end times. ’DAGs triggered but task doesn’t run, possibly due to overlapping schedule and DAG concurrency limit.’
Docker-Based Workflow Orchestration docker, container, compose, run, docker compose, service 613 Problems in Docker Compose environments with service orchestration and path bindings. ’Inside Docker, unable to mount volume correctly, workflow steps fail at execution time.’
Installation and Environment Setup Errors error, install, try, run, apache 611 Environment setup and dependency resolution errors, particularly during installation. ’Error installing Apache Airflow on Windows—command fails at dependency resolution.’
Kubernetes Pod Execution and Management pod, kubernetes, use, run, execute 509 Kubernetes-based execution problems, such as pod lifecycle errors and volume mounting issues. ’Kubernetes pod completes but logs missing, workflow stalled on resource attachment.’
XCom and Cross-Task Data Handling in Airflow xcom, value, task, use, pass, return, operator 335 Cross-task communication failures, especially with return values or Airflow’s XCom. ’XCom returns string list instead of single string, task fails in downstream operator.’
Spark Workflow Submission and Runtime Errors spark, job, submit, use, pyspark 231 Difficulties in running Spark workflows, job submission, and Spark environment setup. ’PySpark task stuck in Spark Standalone Cluster, job never exits execution state.’
Granular Analysis of Workflow Creation and Scheduling: To deepen our understanding of how developers construct workflows, we performed a fine-grained BERTopic analysis on the broader topic of Workflow Creation and Scheduling. This decomposition revealed eight distinct subtopics, detailed in Table 3, covering challenges in DAG syntax, cross-task communication, cloud-based orchestration, and environment configuration. Issues range from syntactic errors in DAG files and misconfigured schedules to complex platform-specific concerns in Spark, Kubernetes, and Docker. This analysis highlights that creating reliable and maintainable workflows requires both domain-specific syntax fluency and robust environment orchestration practices.
Table 4: Granular Subtopics under Distributed Task Management
Subtopic Keywords Post Count Description Representative Example
Task Initialization and Worker Setup task, use, run, try, worker, django, work 3146 Configuring task workers, setting up runtime environments, and customizing task initialization logic. ’Detect and register tasks in Django Python, facing problems with dynamic imports during task loading.’
Framework-Level Integration django, task, run, use, file, error 2463 Challenges in Celery-Django integration, especially related to app setup, module loading, and execution. ’Best way to run an indefinite Celery task within Django app while preserving state?’
Scheduling and Dependency Management task, worker, run, use, chain, queue, execute 1679 Managing dependencies and chaining across distributed task queues and async calls. ’Group result chain stuck in Celery, chained tasks execute out of order.’
Message Broker Access and Authentication rabbitmq, task, queue, worker, message, broker 1077 Connection/authentication issues with message brokers like RabbitMQ and Redis. ’RabbitMQ creates idle queues, no consumer connection, task never received.’
State Management via Redis Queues redis, task, use, run, broker, worker 488 Configuring Redis-based brokers for task routing, storage, and communication reliability. ’Main cause for Redis connection error in Celery when using Django + Redis broker URL?’
Fault Tolerance and Retry Mechanisms retry, fail, exception, handle, task, error, recover 307 Handling retry policies, exceptions, and implementing fault-tolerant execution patterns. ’Handle retry logic after task failure with custom exception handler.’
Worker Deployment and Execution Management celery, task, worker, concurrency, run, queue 287 Deploying Celery workers at scale, managing concurrency, and orchestrating worker pools. ’Run multiple Celery workers in parallel on different queues using Docker.’
@@
The core challenge of this topic arose from the complexity of executing, scheduling, and managing tasks across distributed environments. Developers often struggle with task execution failures, misconfigured workers and queues, and difficulties in periodic task scheduling using SWSs such as Celery. For example, in one post, we found that one developer faced challenges in integrating Celery with Django, especially in configurations and dependency management. Furthermore, scalability and performance bottlenecks emerge as systems grow, requiring careful tuning of workers, queues, and memory allocation. These challenges highlight the need for better tooling, documentation, and automation to streamline distributed task execution in SWSs.
Granular Analysis of Distributed Task Management: To investigate the operational complexities of distributed task systems, we conducted a fine-grained BERTopic analysis on the broader topic of Distributed Task Management. As shown in Table 4, this yielded seven subtopics covering key areas such as worker setup, Celery integration, broker connectivity, and task dependency coordination. Posts reveal persistent difficulties in configuring task runtimes, linking asynchronous jobs, and managing reliability across tools like Celery, RabbitMQ, and Redis. These results highlight the technical overhead associated with scaling and maintaining distributed task workflows, particularly in fault-prone and loosely coupled execution environments.
Table 5: Granular Subtopics under Processor for Data Processing
Subtopic Keywords Post Count Description Representative Example
General Processor Usage in Java Workflows processor, use, file, java, apache, datum, flow 2900 General usage of processors in Java-based workflows, often involving file operations and flow file manipulation. ’Record maximum timestamp value from flow file content and save it.’
NiFi Execution and Error Troubleshooting use, error, try, apache, run, work, docker 828 Challenges in executing and debugging NiFi processors, including setup errors and environment compatibility. ’Make run-nifi.bat work on Windows; getting runtime errors with processor flow.’
Tabular Data Column Configuration column, value, use, table, datum, csv, processor 426 Issues related to modifying or appending tabular data (e.g., columns) in CSV or table-based processors. ’Add column to CSV using predefined values via Apache processors.’
JSON Conversion and Attribute Processing json, use, value, convert, attribute, array, parse 396 Difficulties in converting JSON data, extracting attributes, and handling multiple nested objects. ’Split multiple JSON objects from a single string into flow files using EvaluateJsonPath.’
Kafka Topic and Message Handling kafka, topic, message, use, datum, consumer, offset 187 Integration issues with Kafka, focusing on consuming messages and managing topics and offsets. ’Read Kafka message starting from last committed offset and process it in NiFi.’
Script-Based Data Routing with Conditions route, attribute, match, regex, path, type 166 Conditional routing using processors like RouteOnAttribute, typically based on regex or flowfile metadata. ’Use RouteOnAttribute to separate XML from JSON files before PutFile processor.’
FlowFile Filename and Output Path Issues putfile, filename, directory, output, path, original 92 File output configuration challenges such as naming conventions and output directory setup in PutFile. ’Configure PutFile to save using original filename rather than timestamp.’
@@
3. Processor for Data Processing: SWSs provide a rich set of processors covering a wide range of data integration, transformation, routing, and interaction tasks. Practitioners can design sophisticated data pipelines to meet their specific data processing needs by selecting and configuring the appropriate processors. This topic covers 14.02% of the posts. One example of this topic with about 27K views is ’How to use NiFi ExecuteScript processor with Python?’. ExecuteScript processor allows users to write and execute custom scripts within a data flow. This processor provides flexibility for handling complex processing logic that the built-in processors do not cover. Here, the developer tried to use the ExecuteScript processor with a simple Python script but failed to see the output and asked for help to explain how to use it with an example. Another example of this topic with 12K views is ’How to specify output filename with PutFile processor?’. Here, the problem was that the PutFile processor was saving files using a timestamp, but the expectation was to save them with the original filename. Thus, the developer was seeking a solution. This topic indicates developers’ challenges in understanding data processing processors.
This challenge primarily stemmed from the complexity of configuring and optimizing SWS processors for handling large-scale data workflows. Developers struggle with file management, data flow orchestration, and custom processor debugging, as seen in SO posts. Integration with external systems, data consistency, and scalability under high loads further complicate processing pipelines. Misconfigurations in SWS processors often lead to inefficiencies, requiring better tooling and documentation for smoother data processing workflows.
Granular Analysis of Processor for Data Processing: Our analysis identified seven (Table 5) semantically distinct subtopics, highlighting the diverse processing challenges developers encounter. These range from configuring built-in processors to writing custom scripts. Many posts reflect confusion around general usage, particularly when invoking processors alongside file management tasks in Java-based workflows. Other subtopics point to issues specific to NiFi execution and integration with streaming platforms like Kafka. Notably, concerns such as JSON conversion, attribute handling, and managing FlowFile filenames and output paths reveal the fine-grained complexity of data transformation tasks. Challenges in route configuration and tabular data enrichment further underscore the intricacies of building robust data pipelines in SWSs.
@@
Table 6: Granular Subtopics under Data Structures and Operations
Subtopic Keywords Post Count Description Representative Example
DataFrame Usage with Files and Functions use, dataframe, file, datum, try, error, column 1294 Challenges in applying functions or accessing values in dataframes while handling files or large datasets. ’Apply function to large dataframe; RAM gets overloaded and performance drops.’
Basic DataFrame Manipulations and Errors use, dataframe, try, file, worker, run, code 3182 General usage of dataframes involving basic manipulations, error-prone operations, and try-except patterns. ’Worker throws error shortly after starting; tried using dataframe with large CSV.’
Parquet File Handling in Pipelines parquet, file, read, use, dataframe, write 227 Difficulties in reading and writing Parquet files, especially with empty returns or schema mismatches. ’Read Parquet file returns empty dataframe; trying to debug file structure.’
This challenge occurred primarily due to the complexity of efficiently managing and manipulating large-scale data in distributed computing environments. Developers faced difficulties in optimizing data structures like dataframes and collections across frameworks such as Dask, MongoDB, and Hadoop due to limitations in indexing, partitioning, and parallel processing. Issues include handling large datasets without performance bottlenecks, ensuring seamless operations across distributed systems, and integrating diverse data storage formats. Furthermore, the need for efficient data retrieval, transformation, and storage operations adds another layer of complexity, making it challenging to maintain both performance and scalability.
Granular Analysis of Data Structures and Operations: We decomposed the Data Structures and Operations topic into three coherent subtopics using BERTopic (Table 6) to gain finer insights into developer challenges. The largest subtopic, Basic DataFrame Manipulations, and Errors, includes common operations such as indexing and transformation, often wrapped in error-prone try blocks. DataFrame Usage with Files and Functions captures issues with applying functions to large datasets, frequently leading to performance or memory constraints. Finally, Parquet File Handling in Pipelines reflects difficulties in reading and writing Parquet files, particularly when faced with schema mismatches or empty outputs. These subtopics highlight recurring pain points in structured data management within scientific workflows.
Table 7: Granular Subtopics under Workflow Execution
Subtopic Keywords Post Count Description Representative Example
Execution Job Coordination and Control job, run, workflow, java, use, action, hadoop 777 Challenges in managing execution jobs and controlling task coordination in distributed systems. ’Access local file system not allowed even after setting instructions in Oozie job.’
Workflow Definition and XML Parsing Errors coordinator, workflow, action, job, run, time 406 Errors during XML-based workflow definition or parsing in Hadoop/Oozie systems. ’Difference between workflow, coordinator, and bundle configurations?’
Hive Script and Task Scheduling Issues hive, workflow, file, script, run, error 277 Difficulties with job scheduling, cluster communication, and runtime failures. ’Hive script with input path fails mid-task due to missing temp directories.’
Spark Job Execution and Submission Problems spark, run, job, jar, submit 213 Execution management of actions and triggers in workflow engines like Oozie and Spark. ’Pyspark actions run using Spark 1.6 instead of expected 2.2 in cluster job.’
MapReduce and Environment-Specific Workflow Failures job, run, reduce, map reduce 129 Problems with running workflows inside containerized or secured environments. ’Set up cron job to run three map-reduce pipelines; results inconsistent.’
Failure Recovery and Retry Handling fail, retry, exception, handle, error, step 114 Handling failures, retries, or missed steps in scheduled workflows. ’How to rerun failed step in Oozie without re-triggering the full workflow?’
Shell Script and Job Launch Errors shell, command, job, error, script, exit code 104 Configuration and debugging issues related to launching jobs from shell or command line. ’Shell script exits with code 1; not able to trigger Spark job via command.’
@@
The Workflow Execution challenge primarily stemmed from the complexity of managing and orchestrating distributed workflows in systems like Oozie and Hadoop. Developers struggle with job scheduling, dependency resolution, error handling, and resource allocation issues. Ensuring workflows execute correctly across multiple nodes, handling failures gracefully, and debugging execution errors pose significant challenges. Additionally, misconfigurations, incompatibilities between job components, and inefficiencies in workflow automation further complicate execution, making it difficult to achieve reliable and optimized workflow performance at scale.
Granular Analysis of Workflow Execution: To better understand the challenges developers face during the execution phase of workflows, we applied BERTopic to further decompose the topic Workflow Execution. This fine-grained analysis yielded seven subtopics that reflect distinct pain points ranging from job scheduling and runtime failures to XML parsing errors and environment-specific execution issues. As summarized in Table 7, developers frequently reported problems managing execution jobs, coordinating triggers, and handling resource and configuration mismatches in distributed environments like Hadoop, Oozie, and Spark. These insights suggest that workflow execution remains a technically demanding phase that often requires low-level debugging and system-specific tuning.
Table 8: Granular Subtopics under Managing Rules, Input, Output, and Files
Subtopic Keywords Post Count Description Representative Example
Rule Execution and File Usage Errors file, use, run, rule, error, output, try, snakemake 766 Difficulties in rule definition and execution in Snakemake or similar DSLs. ’Access multiple key config file, question on how to structure rules for proper path resolution.’
Wildcard Pattern and Rule Matching file, rule, input, output, wildcard, use, run, snakemake 599 Challenges in managing wildcard patterns and rule matching logic. ’Output file with mixed wildcard stuck, rule not producing expected target.’
Cluster Job Submission and Path Configurations job, cluster, slurm, submit, run, use, node, resource 141 Issues with configuring input and output paths across tasks and pipelines. ’Trying SGE cluster submission, need help managing resource paths for input and output.’
Conda Environment and File Dependency Issues conda, environment, use, rule, module 94 Handling file dependencies, extensions, and dynamic naming in output rules. ’Problem using Python module from conda env within Snakemake rule.’
Snakefile Errors and Missing Inputs snakefile, error, run, file, try, script, use 91 Debugging problems related to missing inputs, incorrect file paths, or unresolved rule targets. ’Run Snakefile from directory containing it, but output not found, rule fails silently.’
@@
This challenge primarily arose from the complexity of defining, tracking, and executing workflow rules in SWSs like Snakemake and Luigi. Developers struggle with specifying dependencies, handling dynamic file paths, and ensuring that input and output files are correctly mapped across different workflow steps. Issues such as incorrect rule execution, persistent reruns of completed tasks, and difficulties in debugging file-related errors further complicate workflow management. Additionally, ensuring reproducibility while dealing with different file formats and execution environments adds another layer of difficulty, making it challenging to maintain efficient and scalable workflows.
Granular Analysis of Managing Rules, Input, Output, and Files: To further analyze developer difficulties in defining and managing rules in SWSs, we decomposed this topic into five distinct subtopics using BERTopic. As summarized in Table 8, these subtopics reflect common challenges in writing, debugging, and executing rule-based logic, particularly in systems like Snakemake. Developers reported issues ranging from wildcard mismatches and file path errors to misconfigured environments and unresolved rule dependencies. These results highlight the complexity of rule management in file-driven workflow systems and the need for clearer error reporting and automation support.
Table 9: Granular Subtopics under Task Dependencies and Management
Subtopic Keywords Post Count Description Representative Example
Target Files and Dependency Errors use, error, target, file, try, code, run 358 Managing task dependencies, ensuring correct order of execution, and dealing with prerequisites in DAGs. ’Visualize simulation, generic example installed but can’t build target due to missing file dependency.’
Input Definitions and Execution Control use, try, joint, plant, constraint, add 457 Using decorators, directives, or annotations to define task inputs, outputs, and execution contexts. ’Construct simulation in pydrake with manipulation framework, inputs not triggering expected behaviors.’
Luigi Task Linking and Execution Failures task, luigi, run, file, pipeline 234 Debugging task failure or non-execution due to missing dependencies or incorrect conditions. ’Run Luigi task depending on another task, only one executes and the other gets skipped.’
Conditional Task Execution with Pydrake error, python, run, pydrake, import, version 88 Handling conditional logic and parameter passing between tasks in dynamic workflows. ’SystemError on pydrake task run, likely due to incorrect dependency configuration.’
@@
This challenge arose from the complexity of orchestrating and managing task execution in SWSs like Luigi and Drake. Developers face difficulties in defining task dependencies, handling failures, and ensuring that tasks execute in the correct sequence without unintended bottlenecks. Issues such as unregistered task types, state synchronization failures, and debugging interdependent tasks further complicate the process. Additionally, efficiently managing parallel execution while ensuring resource constraints and data consistency presents another layer of difficulty, making it challenging to maintain reliable and scalable task workflows.
Granular Analysis of Task Dependencies and Management: To better capture the nuanced challenges of managing task coordination in workflows, we applied BERTopic to the broader topic Task Dependencies and Management, resulting in four focused subtopics. As shown in Table 9, these subtopics highlight the recurring difficulties developers face when defining task hierarchies, passing parameters, and managing execution dependencies across tools like Luigi and pydrake. Common concerns include task non-execution due to unmet conditions, debugging multi-stage workflows, and integrating conditional logic. These findings underscore the need for clearer dependency resolution mechanisms and improved diagnostic support in task-based workflow engines.
@@
8. Data Transformation: In SWS, this topic discusses converting data from one format or structure to another, enabling easier analysis, visualization, or integration with other datasets. This involves altering the shape or appearance of data objects while preserving their content and, for instance, using libraries like Jolt to transform JSON data from one format to another. This topic covers around 1.26% of posts of our SO dataset. One example of this topic, with around 10K views, is titled ’Is it possible to concatenate the values of JSON attributes using JOLT?’. In this instance, the developer inquired whether JOLT could be used to concatenate the values of three JSON attributes into a new attribute and also asked about potential alternative approaches.
The Data Transformation challenge originated from the complexity of structuring, mapping, and converting data formats across different processing frameworks. Developers struggle with transforming nested JSON structures, applying transformation rules using tools like Jolt and NiFi, and ensuring data integrity during the conversion process. Issues such as handling dynamic key-value mappings, concatenating attributes, and efficiently parsing large data sets add to the difficulty. Furthermore, inconsistencies in schema definitions, missing values, and compatibility issues between input and output formats make it challenging to implement seamless data transformation workflows.